### v2

# DynamicKV: Task-Aware Adaptive KV Cache Compression for Long Context LLMs

Abstract
Efficient KV cache management in LLMs is crucial for long-context tasks like RAG and summarization. Existing KV cache compression methods enforce a fixed pattern, neglecting task-specific characteristics and reducing the retention of essential information. However, we observe distinct activation patterns across layers in various tasks, highlighting the need for adaptive strategies tailored to each task’s unique demands. Based on this insight, we propose DynamicKV, a method that dynamically optimizes token retention by adjusting the number of tokens retained at each layer to adapt to the specific task. DynamicKV establishes global and per-layer maximum KV cache budgets, temporarily retaining the maximum budget for the current layer, and periodically updating the KV cache sizes of all preceding layers during inference. Our method retains only $1.7\%$ of the KV cache size while achieving $\sim 85\%$ of the Full KV cache performance on LongBench. Notably, even under extreme compression ($0.9\%$), DynamicKV surpasses state-of-the-art (SOTA) methods by 11% in the Needle-in-a-Haystack test using Mistral-7B-Instruct-v0.2. The code will be released.
DynamicKV: Task-Aware Adaptive KV Cache Compression for
Long Context LLMs
Xiabin Zhou, Wenbin Wang, Minyan Zeng, Jiaxian Guo, Xuebo Liu, Li Shen, Min Zhang, Liang Ding {xiabinzhou0625, liangding.liam}@gmail.com
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Radford, 2018; Radford et al., 2019) are exerting a considerable influence in the field of natural language processing (NLP), driving advancements in summarization, translation, code generation, etc. (Chiang et al., 2023; Zhong et al., 2023; Peng et al., 2023; Lu et al., 2024; Wang et al., 2024). Recent developments in LLMs (Liu et al., 2024b) have been scaled up to handle long contexts, with LlaMA3 (Dubey et al., 2024) processing up to 32K tokens and InternLM (Cai et al., 2024) handling 1M tokens. However, scaling LLMs to handle extended contexts inherently incurs a substantial delay due to the quadratic complexity of attention mechanisms with increasing context length. A widely adopted solution to alleviate these delays is caching the key and value (KV) states of previous tokens (Waddington et al., 2013). Despite this optimization, handling long sequences still demands substantial memory (e.g., maintaining a KV cache for 100K tokens in LlaMA2-7B (Touvron et al., 2023) consumes over 50 GB of memory).
To address this issue, recent studies have explored the optimization of KV caching, including KV cache quantization (Kang et al., 2024; Hooper et al., 2024), token dropping (Zhang et al., 2024b; Xiao et al., 2023), architectural improvements to Transformers (Sun et al., 2024), KV cache fusion (Nawrot et al., 2024), and hierarchical sharing and constraints(Liu et al., 2024a; Brandon et al., 2024). Existing KV cache compression methods enforce a fixed pattern (as shown in Figure 1), such as a hierarchical pyramid structure (Zhang et al., 2024a) or a structure similar to FastGen’s fixed internal pattern (Ge et al., 2023), or they fix the length of the KV cache to selectively retain tokens across different layers (Zhang et al., 2024b; Li et al., 2024). However, LLMs require different numbers of layers when handling different types of tasks. For example, for knowledge-based question-answering tasks, only the first few layers are needed to achieve high accuracy, while for complex reasoning tasks (e.g., mathematics and code generation), more layers are often required to achieve higher accuracy (Elhoushi et al., 2024). Thus, we raise a question: Do different types of tasks all follow a fixed pattern?
Figure 1: Comparison of DynamicKV with traditional methods in maintaining KV cache size across layers. Left: the structure difference: (a) Retain all KV cache. (b) Fixed KV cache for each layer (e.g., StreamingLLM, H2O, SnapKV). (c) Hierarchically decreasing pyramid KV cache retention. (d) Ours DynamicKV: layer-aware adaptive KV cache retention. Right: average accuracy on different KV cache retention.
To examine this question, we aim to systematically investigate the design principles of the KV cache compression across different tasks. Inspired by Zhang et al. (2024a), we first investigate how information flow is aggregated through attention mechanisms across different layers in four types of tasks, including single- and multi-document QA, summarization, synthetic tasks and code completion. We find that the attention distribution varies for different types of tasks. For example, in summarization tasks, the upper layers require a small KV cache size, while code completion tasks need larger KV cache sizes in the upper layers. This implies that for code completion tasks, upper layers require maintaining a larger KV cache size, in contrast to PyramidKV (Zhang et al., 2024a), where the KV cache size decreases as the layer depth increases.
Building on this insight, we propose a task-aware adaptive KV cache compression method, named DynamicKV. Specifically, we first calculate an attention score for the most recent few tokens and all other tokens, which in RAG (Lewis et al., 2020) can be viewed as calculating the relevance of the most recent query to the retrieved text. Then, we preset a temporary storage to hold the temporary KV cache states and gradually calculate the size of the final retained temporary storage at each k layer by calculating the size of the correlation mean. It should be noted that at each update, the value is gradually normalized, and the retained temporary storage at each layer is always smaller than the previous one. This temporary storage is determined by the number of tokens that need to be retained, and its size is much smaller than the original cache, thus imposing minimal memory overhead.
We validate our DynamicKV on 16 datasets from LongBench (Bai et al., 2023), demonstrating robust performance across multiple models, including LlaMA-3-8B-Instruct (Dubey et al., 2024), Qwen-2-7B-Instruct (Yang et al., 2024), Mistral-7b-Instruct-v0.2 (Jiang et al., 2023), InternLM-2.5-7b-Chat-1M (Cai et al., 2024). Our DynamicKV exhibits superior overall effectiveness compared to conventional fixed-pattern methods (Zhang et al., 2024b; Li et al., 2024; Nawrot et al., 2024). Notably, DynamicKV can retain full performance while utilizing only 6.9% of the tokens, and in extreme scenarios, it preserves 90% of the performance with just 1.7% of the tokens. Furthermore, experiments on the Needle in a Haystack benchmark revealed that DynamicKV significantly outperforms state-of-the-art (SOTA) methods.
2 Related Work
Potential patterns of attention in LLMs. The Transformer architecture (Vaswani, 2017) becomes a cornerstone in NLP by stacking multiple layers to progressively refine input data. BERT (Devlin, 2018), a model based on this architecture, Jawahar et al. (2019) demonstrates that intermediate layers encode a rich hierarchy of linguistic information: from surface-level features at the bottom, through syntactic features in the middle, to semantic features at the top. This indicates that models are capable not only of understanding lexical information but also of grasping more complex linguistic structures.
For decoder-only LLMs, Fan et al. (2024) observes that not all layers are necessary for simple tasks, as intermediate layers can often achieve comparable performance to the final layer. Techniques like (Elhoushi et al., 2024), which involve increasing dropout in lower layers during training, allow the model to exit computation early, reducing resource consumption.
Figure 2: Average token retention across layers in LlaMA for different tasks, including (a) Single/Multi-Document QA, (b) Summarization, (c) Synthetic Task, and (d) Code Completion. There is a sharp decrease in token retention after the first layer, followed by varying patterns of fluctuation. Peaks are observed around Layer 15 and towards the final layers.
To optimize model inference efficiency, especially in terms of KV cache compression, Brandon et al. (2024) proposes cross-layer attention(CLA), which can reduce the KV cache size by at least half by sharing cross-layer attention, significantly lowering memory usage. Ada-KV (Feng et al., 2024) visualizes attention distributions across all layers have also shown that attention patterns dynamically evolve as the layers progress. Inspired by these findings, we aim to dynamically select and adjust the number of tokens to retain per layer, combining inter-layer redundancy identification with efficient KV cache management. This approach aims to maintain high-quality output while improving inference efficiency.
Token drop strategies in KV cache compression.
Token drop is a strategy designed to reduce memory usage by selectively retaining the most influential tokens in the KV cache during the inference phase of LLMs. Due to its plug-and-play nature, the token drop method can often be applied to different models without incurring any additional costs. FastGen (Ge et al., 2023) evicts unnecessary contexts and discards non-special tokens based on the recognized structure of attention modules by effectively analyzing the token information within attention patterns. Scissorhands (Liu et al., 2024c) exploits the hypothesis of the persistence of importance, suggesting that tokens with significant influence at one point will continue to impact future generations. By using attention scores as a metric and applying a Least Recently Used (LRU) cache eviction strategy, it discards non-critical tokens to optimize memory usage. StreamingLLM (Xiao et al., 2023) leverages the characteristics of attention sinks in LLMs to focus on streaming processing with dynamic adjustment of the KV cache. H2O (Zhang et al., 2024b) proposes a scoring function based on accumulated attention scores for greedily evicting KV pairs during generation. SnapKV (Li et al., 2024) primarily achieves compression by selectively targeting key positions for each attention head. PyramidKV (Zhang et al., 2024a) identified the phenomenon of massive activation and adopted a hierarchical structure to optimize the number of KV cache entries retained at each layer. Although the PyramidKV approach considers the varying information density across different layers, its pyramidal pattern does not generalize across multiple models or tasks. LazyLLM (Fu et al., 2024) utilizes dynamic token pruning and an Aux Cache mechanism, allowing the model to select different subsets of tokens from the context at various generation steps, even reviving tokens pruned in previous steps. Ada-KV (Feng et al., 2024) breaks from the conventional approach of uniform budget allocation across attention heads within layers, and optimizes the eviction loss upper bound, leading to improved performance under various memory constraints when integrated with SnapKV and PyramidKV.
3 Observation
To systematically investigate the attention mechanism across layers in LLMs for long-context inputs, we conduct a fine-grained analysis of four tasks: single- and multi-document question answering (QA), summarization, synthetic tasks, and code completion. The main target is to investigate the distribution of attention in these various tasks, thereby enhancing our understanding of how the model aggregates dispread information within long-context inputs to generate accurate responses.
In particular, we focus our analysis on LlaMA (Dubey et al., 2024), visualizing the distribution and behavior of attention across layers to gain deeper insights into its internal mechanisms. Inspired by Zhang et al. (2024a), we calculate the average attention scores between the most recent tokens and all other tokens. Based on these scores, we then identify the top-k (128 multiplied by the number of layers) tokens with the highest attention across all layers, resulting in a layer distribution map denoted as Figure 2.
We observe a significant drop in the KV cache size requirement at the lower layers across the four tasks, indicating that only a small KV cache is needed in these layers. In contrast, the upper layers show a clear upward trend, suggesting that larger KV cache sizes are necessary, particularly in the code completion task, where complex reasoning is required. This phenomenon underscores that tasks involving complex reasoning demand larger KV cache sizes in the upper layers.
4 DynamicKV
During inference, the quadratic complexity of attention calculation results in a significant computational and memory burden, especially when processing long contexts. DynamicKV addresses this issue by focusing on inter-layer attention in large language models (LLMs), determining the appropriate size of KV cache to retain per layer through efficient awareness of inter-layer attention.
Rather than relying on a fixed retention pattern, such as pyramid-shape or average retention all layers, DynamicKV employs a progressive algorithm that dynamically adjusts token retention during the prefill phase. This dynamic retention strategy accelerates the decoding stage while maintaining minimal impact on overall memory usage.
Specifically, we first define layer $l\in\displaystyle\mathbb{R}^{L}$ and head $h\in\displaystyle\mathbb{R}^{H}$ in LLMs. For the calculation of attention scores, we use weights $W_{Q}\in\displaystyle\mathbb{R}^{N\times N}$, $W_{K}\in\displaystyle\mathbb{R}^{N\times N}$, and $W_{V}\in\displaystyle\mathbb{R}^{N\times N}$, with the input query embedding denoted as $X\in\displaystyle\mathbb{R}^{N\times M}$, $N$ is the dimension of the hidden size, and $M$ is the length of input tokens. Traditional token drop methods often consider the most recent tokens as the important ones for producing output information, as they retain relevant information needed for generating answers. We refer to these tokens collectively as the current window, with the window size denoted as $ws$. In the prefill phase, we adopt the method from Li et al. (2024), Zhang et al. (2024a), where the attention score is calculated by averaging over the current window and previous tokens, followed by pooling. The formula is as follows:
$A_{l,h}=\text{pooling}(\frac{1}{ws}\sum_{i=1}^{ws}\text{Attention}(X_{i},W_{Q},W_{K})),$ (1)
where pooling helps better understand the context and $A_{l,h}$ denotes the attention score for the l-th layer and the h-th head. This approach allows us to effectively pool the attention scores, ensuring that key tokens are retained based on their relevance to both the current window and previous context.
Figure 3: Overview of our DynamicKV structure and KV cache compression comparison. Left: Layer-wise KV cache retention mechanism in transformer architectures. Right: Our proposed DynamicKV framework employs stage-wise dynamic updating to maintain KV cache within predefined memory budgets, with task-specific visualization showing KV cache preservation patterns across layers.
Next, we set a fixed retention budget. Specifically, to ensure a fair comparison with other methods, we introduce the average retention length per layer, denoted as $wt$, and a scaling ratio, $r_{max}$. The calculation formula is as follows:
$bs=(wt-ws)\times r_{max},$ (2)
where $bs$ represents the size of retained KV cache across all layers. Next, we design a layer-aware progressive dynamic KV cache compression method. The prefill phase of LLMs involves a hierarchical forward process, where for each layer, we retain a KV cache of length $bs$ when computing $A$. Additionally, every $m$ layers, we perform an update across the current and all previous layers. Specifically, for each layer, we use a top-$k$ strategy to retain the largest $bs$ values from $A_{l}$, where $A_{l}$ represents the attention scores of layer $l$. The formula for this process is as follows:
$A_{l}^{\prime}=\text{TopK}(A_{l},bs).$ (3)
Next, we extract the indices in the original $A_{l}$ that correspond to the values in $A_{l}^{\prime}$. The KV cache at these indices is retained as the compressed KV cache. Specifically, the retained KV cache is defined as:
$\text{KV}_{l}^{\prime}=\text{KV}_{l}[A_{l}^{\prime}.\text{indices}],$ (4)
where $A_{l}^{\prime}.\text{indices}$ represents the indices of the top-$k$ values in $A_{l}$. This ensures that the KV cache is compressed efficiently, retaining only the most important tokens for each layer while minimizing memory usage.
To ensure that the memory required for hierarchical transmission remains small, the KV cache of each layer is initially compressed as described above. For every $m$ layer, we extract $A$ and perform a unified normalization across the completed layers, updating them layer by layer to ensure consistency across the entire hierarchy.
First, we fix the final size of the KV cache to be retained, which is calculated as $(wt-ws)\times H\times l$, where $H$ is the number of heads and $l$ is the number of layers. Then, for each layer, the attention score $A$ is used to compute the length to retain for each layer $C_{l}$ via a top-$k$ strategy. The retention lengths for the first $m$ layers are then normalized to obtain a budget length $Z$, ensuring that the retention is distributed effectively across layers. The specific formula is as follows:
[…]
$C_{l}=\text{Norm}(\text{Count\_Elements}\left(I\right))$ (6)
$Z=\left[\frac{bs\times t}{\textit{max}(C_{l})}\ ~{}\textit{for}\ t\ \in\ C_{l}\right]$ (7)
$r=\frac{\sum Z}{(wt-ws)\times L},Z=\left[\frac{k}{r}\ ~{}\textit{for}\ k\ \in\ Z\right]$ (8)
The KV cache is further updated layer by layer based on this normalized budget, progressively refining the retained information to align with the overall compression strategy. The above process can be expressed as Algorithm 1.
5 Experiments
We conduct comprehensive comparative and ablation experiments to verify the effectiveness of our DynamicKV. In Section 5.1, we introduce the models, datasets and baselines used in our experiments. Section 5.2 provides a performance comparison between DynamicKV and baseline approaches. Next, in Section 5.3, we present the results of DynamicKV on the Needle in Haystack Task. Finally, in Section 5.4, we conduct an ablation study on the parameters of our method to validate its feasibility.
5.1 Implementation details
Models and Context Length. We utilize the official checkpoints of recently released models from huggingface including LlaMA-3-8B-Instruct (Dubey et al., 2024), Qwen-2-7B-Instruct (Yang et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and InternLM-2.5-7B-Chat-1M (Cai et al., 2024) as our base models, which support context lengths of 8k, 32k, 32k, and 1M tokens respectively.
Datasets.
LongBench is a comprehensive benchmark for evaluating the contextual understanding capabilities of LLMs. For our comparative experiments, we use 16 English datasets from this benchmark, specifically NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), MultiFieldQA-en, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), MuSiQue (Trivedi et al., 2022), GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), MultiNews (Fabbri et al., 2019), TREC (Li and Roth, 2002), TriviaQA (Joshi et al., 2017), SAMSum (Gliwa et al., 2019), PassageCount, PassageRetrieval-en, LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023). These cover key long context application scenarios such as Single-Document QA, Multi-Document QA, Summarization, Few-shot Learning, Synthetic Tasks, and Code Completion. Additionally, for the experiment on the Needle in Haystack task, we test the models across their maximum length ranges [8k, 32k, 1M] using the PaulGrahamEssays dataset.
Baselines.
We evaluate the recent fixed-pattern token-dropping methods, including: (1) StreamingLLM, which utilizes attention sinks and rolling KV caches to retain the most recent tokens. (2) H2O, which employs a Heavy Hitter Oracle for KV cache eviction. (3) SnapKV, which selects important tokens for each attention head through clustering. (4) PyramidKV, which introduces a pyramid pattern where layers select important tokens in a monotonically decreasing manner.
5.2 Comparative experiments on LongBench
Model Size Method Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg.
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
18409 3619 4559 9151 4887 11214 8734 10614 2113 5177 8209 6258 11141 9289 1235 4206
LlaMA-3-8B -Instruct FullKV 25.16 31.81 39.59 43.09 36.15 21.77 28.62 23.34 26.33 75.00 90.50 42.36 5.20 69.25 59.04 53.93 41.95
128 StreamingLLM 17.85 9.50 23.09 37.84 29.02 16.77 17.91 20.42 20.16 44.00 73.00 30.00 5.80 69.50 48.38 49.31 32.03
H2O 21.58 12.54 28.49 37.13 32.36 18.88 20.23 22.16 21.14 39.00 86.62 39.19 5.50 69.50 57.39 54.46 35.39
SnapKV 21.71 12.37 32.38 37.44 30.48 19.50 19.06 21.36 20.07 45.5 87.74 38.15 5.50 68.85 57.42 54.61 35.76
PyramidKV 22.26 16.65 30.73 38.97 29.28 19.19 19.92 22.06 20.87 68.00 88.95 38.23 5.92 69.50 57.20 51.54 37.45
ours 22.10 14.93 32.94 41.06 27.98 21.18 20.03 22.06 21.28 65.50 89.61 38.70 5.13 69.50 58.01 54.00 37.75
512 StreamingLLM 19.03 12.78 28.67 37.83 29.97 16.55 20.30 20.94 24.56 61.00 75.43 30.82 5.86 69.50 51.93 49.98 34.70
H2O 22.84 16.80 32.36 41.43 34.07 19.30 22.28 22.81 23.69 41.00 90.46 40.19 5.54 69.50 57.52 55.43 37.20
SnapKV 24.62 22.78 37.88 42.96 34.82 20.65 22.63 22.54 23.93 70.00 90.39 40.30 5.74 69.50 60.27 55.85 40.30
PyramidKV 24.48 23.51 36.14 42.33 31.95 20.73 23.37 23.01 24.37 72.50 90.43 40.54 5.88 69.50 59.25 54.87 40.18
ours 24.78 24.76 36.84 44.13 33.25 20.82 23.00 22.76 24.14 72.50 90.39 40.76 5.78 69.50 61.40 56.91 40.73
Mistral-7B -Instruct-v0.2 FullKV 26.63 32.99 49.34 42.77 27.35 18.77 32.87 24.24 27.10 71.00 86.23 42.96 2.75 86.98 56.93 54.49 42.71
128 StreamingLLM 16.58 14.76 30.36 28.13 21.76 11.98 18.26 19.02 19.16 43.50 74.12 28.50 2.50 31.81 43.65 41.19 27.83
H2O 21.66 21.64 38.60 30.96 20.63 13.02 20.65 22.61 22.08 39.00 82.19 39.75 3.16 79.98 51.25 48.20 34.71
SnapKV 20.11 21.28 42.98 37.51 22.31 14.43 19.19 21.89 21.01 48.00 83.77 40.44 2.51 66.99 51.64 48.57 35.16
PyramidKV 22.11 22.52 43.04 33.57 22.98 15.69 20.56 22.52 21.36 65.50 83.84 40.03 2.89 67.26 51.51 46.42 36.36
ours 22.05 23.65 43.08 36.03 22.60 15.23 21.35 23.11 22.19 68.00 84.79 41.02 4.20 70.11 52.45 47.41 37.33
512 StreamingLLM 19.05 17.21 36.82 30.64 21.84 10.56 24.47 19.84 25.48 62.00 72.82 29.49 2.71 19.25 46.15 42.55 30.06
H2O 22.33 25.75 44.09 32.76 22.88 14.96 23.53 22.96 24.53 41.50 85.53 41.54 3.39 86.20 55.11 50.81 37.37
SnapKV 24.95 27.97 49.04 39.93 25.18 17.64 24.14 23.69 24.47 67.50 86.04 41.14 2.90 86.98 56.73 53.11 40.71
PyramidKV 23.49 28.79 48.71 41.00 25.64 16.35 24.79 23.52 24.49 69.50 86.20 42.58 3.53 81.81 55.45 51.67 40.47
ours 25.63 29.11 48.41 39.85 26.62 16.72 24.73 23.72 24.83 70.50 86.74 43.01 3.20 83.57 55.40 52.35 40.90
Qwen2-7B -Instruct FullKV 25.14 42.35 45.04 14.80 14.13 9.23 36.35 23.79 26.51 76.50 89.16 45.23 6.50 75.50 60.30 60.78 40.71
128 StreamingLLM 19.25 23.63 26.51 14.00 15.30 7.46 18.07 19.30 18.30 47.00 77.92 31.57 6.50 17.00 42.52 41.94 26.64
H2O 20.33 30.43 34.22 13.61 13.37 7.81 20.72 21.66 18.44 40.00 86.94 42.17 7.00 70.50 53.45 53.76 33.40
SnapKV 22.26 31.62 38.95 16.05 17.71 7.66 18.91 21.41 18.21 46.00 87.61 42.01 6.50 63.50 54.87 53.03 34.14
PyramidKV 20.50 31.70 39.95 18.54 18.54 8.85 19.24 20.47 18.18 60.00 87.98 39.71 7.00 49.00 48.77 47.91 33.52
ours 22.77 35.57 42.62 14.80 16.35 8.31 21.41 21.97 19.56 58.00 88.18 40.93 6.50 70.00 53.58 52.50 35.82
512 StreamingLLM 20.47 26.97 32.64 14.31 14.39 6.82 25.70 19.31 24.88 66.00 76.56 32.11 8.00 15.50 46.58 44.20 29.65
H2O 22.88 34.28 41.40 13.30 14.60 8.31 23.69 22.07 22.72 39.50 88.75 43.91 6.00 72.00 58.83 57.83 35.63
SnapKV 23.86 38.61 44.65 15.60 14.62 9.13 24.56 22.39 23.07 70.00 89.31 43.32 5.00 72.00 58.67 60.74 38.47
PyramidKV 24.47 37.60 43.51 14.48 12.83 8.99 23.59 22.30 22.41 74.00 89.21 43.40 6.50 74.00 57.67 56.14 38.19
ours 24.66 40.44 45.30 15.42 13.89 8.46 25.51 22.77 22.92 74.00 89.27 43.18 7.00 74.00 60.38 59.33 39.16
InternLM-2.5-7B -Chat-1M FullKV 22.42 27.61 39.98 40.92 33.48 26.68 33.01 25.18 26.28 72.50 86.76 39.76 2.91 100.00 55.86 57.95 43.21
128 StreamingLLM 17.91 13.02 24.31 24.27 16.01 11.29 17.29 20.62 18.06 48.5 67.53 21.93 0.82 87.39 43.45 42.79 29.70
H2O 16.16 17.71 27.94 26.83 17.83 17.81 13.99 22.59 16.9 39.50 81.87 32.15 1.32 96.50 48.30 47.27 32.79
SnapKV 19.65 17.44 35.29 27.36 18.58 19.79 12.76 22.42 16.31 48.00 80.23 31.35 0.95 95.00 49.47 48.22 33.93
PyramidKV 18.80 17.35 33.48 31.16 20.05 19.02 14.65 22.02 17.40 69.50 80.87 32.02 1.23 95.00 47.13 44.73 35.28
ours 17.93 19.89 34.15 31.50 19.03 20.60 15.14 22.41 18.15 70.00 83.09 32.44 0.86 95.50 49.33 47.16 36.07
512 StreamingLLM 17.58 15.86 26.55 26.68 16.69 11.01 25.96 21.33 25.57 65.00 67.16 21.71 0.95 87.56 43.58 42.76 32.25
H2O 15.33 19.84 32.41 27.88 20.10 21.13 16.91 22.99 21.49 41.00 84.38 34.76 1.23 96.50 48.46 50.00 34.65
SnapKV 16.86 23.28 36.24 32.14 19.89 23.21 17.69 23.18 22.44 71.00 84.05 34.34 1.00 96.50 50.32 53.34 37.84
PyramidKV 17.62 21.08 37.52 32.21 21.31 22.03 19.37 24.06 22.22 73.00 83.94 34.61 1.05 95.50 50.45 49.72 37.86
ours 17.77 23.87 37.74 32.98 21.13 20.85 19.13 23.49 22.48 75.00 84.89 36.70 0.91 95.50 50.70 51.08 38.39
Table 1: Performance comparison on the LongBench dataset for full KV cache, previous methods (StreamingLLM, H2O, SnapKV, PyramidKV), and our DynamicKV method, with KV cache sizes of 128 and 512, using models including LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.2, QWen2-7B-Instruct, and InternLM-2.5-Chat-1M. Bold indicates the best performance.
With the total KV cache size fixed at 128 and 512, we compare the performance retention of StreamingLLM, H2O, SnapKV, PyramidKV, and our proposed method, DynamicKV, relative to FullKV. As shown in Table 1, DynamicKV demonstrates stable improvements even while maintaining an extremely low KV cache size relative to the total context (128: 1.7%; 512: 6.9%). Specifically, with the cache size of 128, DynamicKV outperforms the best alternative by 0.3%, 0.97%, 1.68%, and 0.79% on LLaMA, Mistral, Qwen, and InternLM, respectively, retaining 90%, 87%, 78%, and 83% of the overall performance. Moreover, with a cache size of 512, DynamicKV surpasses the highest-performing method by 0.43%, 0.19%, 0.69%, and 0.53% on the same models, retaining 97%, 96%, 96%, and 89% of FullKV’s performance. The data in the table clearly demonstrate DynamicKV’s effectiveness under extreme compression, achieving nearly FullKV-level performance with just 6.9% of the cache size. The experimental results show that DynamicKV can improve the effect of complex tasks such as code completion more obviously based on maintaining PyramidKV performance, and greatly improve the performance upper limit of lower KV cache size.
5.3 Visualization on Needle-in-Haystack Task
(a) FullKV
(b) StreamingLLM
(c) PyramidKV
(d) DynamicKV
Figure 4: Performance Comparison on the Needle in a Haystack Task Using Mistral-7B-Instruct-v0.2.
The needle-in-a-haystack test involves inserting key information at random positions within a long context and setting answers to evaluate whether LLMs can accurately detect critical information in extensive contexts. To further illustrate the effectiveness of our approach in compressing the KV cache, we conduct additional experiments using Mistral on the needle-in-a-haystack task, focusing on maintaining an optimal size for the KV cache.
As shown in Figure 4, we insert information at various positions in the Paul Graham Essays dataset and extract answers by prompting the model to generate responses. The green blocks indicate that the response matches the contents of the needle, but the colour change from yellow to red indicates that the response is more irrelevant to the needle.
We test a fixed KV cache size of 64 using FullKV, StreamingLLM, H2O, SnapKV, PyramidKV, and the DynamicKV method. The results indicate that DynamicKV maintains 90% of the model’s performance even under extreme compression, improving accuracy by 57%, 37%, 41%, and 11% compared to the other methods, respectively. Additionally, the figure shows that with a context length of up to 7000, the extreme compression of DynamicKV nearly achieves full scores, and even beyond 7000, it shows significant improvements compared to other approaches. This finding illustrates that DynamicKV has a distinct advantage in hierarchical token selection and confirms that the number of critical tokens contained at different layers is always dynamic.
5.4 Ablation Study
KV size LlaMA-3-8B- Instruct Mistral-7B- Instruct-v0.2 Qwen2-7B- Instruct InternLM2.5-7B- Chat-1M
64 34.93 33.95 32.67 33.67
96 36.70 36.22 34.85 35.31
128 37.75 37.33 35.82 36.07
256 39.83 39.23 36.98 37.29
512 40.73 40.90 39.16 38.39
1024 41.22 41.48 39.72 38.86
Table 2: Performance of DynamicKV with different KV cache size.
In this study, we investigate the performance of the DynamicKV mechanism across varying key-value cache sizes. The results, as shown in Table 2, reveal a consistent improvement in performance with an increase in the cache size for all evaluated models. For the LlaMA-3-8B-Instruct, the performance metric improved from 34.93 to 41.22 as the key-value cache size was increased from 64 to 1024. This improvement is also applicable to other models. These findings underscore the effectiveness of the DynamicKV cache in leveraging KV cache compression to maintain the capabilities of long context. Notably, a larger cache capacity is generally associated with superior performance. Nonetheless, it is essential to strike a balance when selecting the cache size, taking into account the practical constraints related to storage and computational resources.
☞ More analyses in Appendix
In addition to the above discussions, we conduct more related analyses and show them in Appendix: (1) DynamicKV is effectively scalable to other models on the Needle-in-Haystack task (Appendix A.3); (2) Enhancing generation efficiency with 22.5-129% higher token throughput compared to fullKV and superior memory optimization saving up to 11.2% GPU memory while maintaining stable performance (Appendix A.4); (3) Task-Aware adaptability that dynamically adjusts token processing across transformer layers based on input complexity(Appendix A.5). Please refer to the Appendix for more details.
6 Conclusion
In this study, we analyze the intrinsic patterns exhibited by large language models (LLMs) when processing long-context inputs across different task types. Our empirical findings reveal significant variations in the distribution of attention across these task types. Based on this observation, we introduce DynamicKV, a novel layer-aware KV cache compression approach that dynamically adjusts the KV cache size across layers. We evaluate the effectiveness and generalizability of DynamicKV through experiments on 16 datasets from the LongBench benchmark, demonstrating its broad applicability and performance benefits. From the results, we mainly conclude that: (1) a wave-like pattern is followed in complex reasoning tasks (e.g., code completion tasks); (2) a pyramid-like pattern is followed in Synthetic and Summarization tasks; (3) The dynamic hierarchical adaptive DynamicKV approach is capable of formulating a relatively appropriate KV cache retention strategy in accordance with diverse tasks. Particularly, in the circumstance of maintaining an extremely small KV cache size, the effect is significantly enhanced.; In the future, we hope that there is a more suitable method to perform KV cache compression without increasing the computation.
Limitations
Our work has several potential limitations. First, given the limited computational budget, we only validate our DynamicKV on models Scaling up to super-large model sizes (e.g., 70B), and applying DynamicKV to more cutting-edge model architectures will be more convincing model architectures. Second, although we have conducted experiments on multiple tasks including single- and multi-document QA, summarization, synthetic tasks, and code completion, the generalization ability of DynamicKV to other tasks or datasets has not been fully explored. Future work will focus on expanding the application scope of DynamicKV to more diverse tasks and datasets.
Appendix A Appendix
A.1 Model Details
All the model structures and details in our experiment are shown in Table 3.
Configuration LlaMA-3-8B- Instruct Mistral-7B- Instruct-v0.2 Qwen2-7B- Instruct InternLM2.5-7B- Chat-1M
Hidden Size 4,096 4,096 3,584 4096
# Layers 32 32 28 32
# Query Heads 32 32 28 32
# KV Heads 8 8 4 8
Head Size 128 128 128 128
Intermediate Size 14,336 14,336 18,944 14336
Embedding False False False False
Vocabulary Size 128,256 32,000 151,646 92,544
Table 3: Configuration of Models.
Dataset Source Avg length Metric Language #data
Single-Document QA
NarrativeQA Literature, Film 18,409 F1 English 200
Qasper Science 3,619 F1 English 200
MultiFieldQA-en Multi-field 4,559 F1 English 150
Multi-Document QA
HotpotQA Wikipedia 9,151 F1 English 200
2WikiMultihopQA Wikipedia 4,887 F1 English 200
MuSiQue Wikipedia 11,214 F1 English 200
Summarization
GovReport Government report 8,734 Rouge-L English 200
QMSum Meeting 10,614 Rouge-L English 200
MultiNews News 2,113 Rouge-L English 200
Few-shot Learning
TREC Web question 5,177 Accuracy (CLS) English 200
TriviaQA Wikipedia, Web 8,209 F1 English 200
SAMSum Dialogue 6,258 Rouge-L English 200
Synthetic Task
PassageCount Wikipedia 11,141 Accuracy (EM) English 200
PassageRetrieval-en Wikipedia 9,289 Accuracy (EM) English 200
Code Completion
LCC Github 1,235 Edit Sim Python/C#/Java 500
RepoBench-P Github repository 4,206 Edit Sim Python/Java 500
Table 4: An overview of the dataset statistics in LongBench.
A.2 Dataset Details
The data sources, average length, evaluation metrics, language, and data volume of the LongBench(Bai et al., 2023) dataset’s subdatasets are shown in Table 4.
Model StreamingLLM H2O SnapKV PyramidKV DynamicKV
LlaMA-3-8B-Instruct 0.29 0.46 0.80 0.89 0.9
Qwen-2-7B-Instruct 0.22 0.41 0.84 0.86 0.87
Table 5: Configuration of Models.
A.3 Need in a HayStack
The experimental results are presented in Table 5, which illustrates the performance of our DynamicKV approach across various models, including LlaMA and Qwen, in Needle in a Haystack task.
Input Len Output Len Method TTFT (s) TPOT (tok/s) Latency (s) Memory (MB)
8k 2k FullKV 0.66 27.63 74.79 20055
8k 2k Dynamickv 0.70 33.85 61.21 19417
16k 4k FullKV 1.45 19.55 209.56 23859
16k 4k Dynamickv 1.49 33.02 125.52 22051
32k 8k FullKV 3.52 11.65 706.56 31213
32k 8k Dynamickv 3.58 26.69 310.56 27713
Table 6: Efficiency Comparison Between FullKV and DynamicKV
A.4 Efficiency Experiments
We evaluate the efficiency of DynamicKV against the standard method (FullKV) under varying input/output lengths. All experiments are conducted with a fixed context window ($m=128$), measuring Time-to-First-Token (TTFT), Time-Per-Output-Token (TPOT), end-to-end latency, and GPU memory usage. The results are summarized in Table 6.
Key observations include:
  • Short Sequences (8k/2k): DynamicKV improves TPOT by 22.5% (27.63→33.85 tok/s) while slightly increasing TTFT by 6% (0.66s→0.70s), achieving 18.2% lower total latency (74.79s→61.21s) with 638MB memory reduction.
  • Long Sequences (32k/8k): The advantages amplify significantly, with DynamicKV delivering 129% higher TPOT (11.65→26.69 tok/s), 56% lower latency (706.56s→310.56s), and 11.2% memory savings (31213MB→27713MB).
  • Scalability: FullKV shows superlinear TPOT degradation (11.65 tok/s at 32k inputs), while DynamicKV maintains stable throughput through on-demand computation, demonstrating better adaptability to long-context generation.
The experiments demonstrate that dynamic KV caching trades marginal initial latency for substantially better sustained generation speed and memory efficiency, particularly beneficial for long-text generation tasks (>2k output tokens).
Dataset Layer-wise ( 1 - 32 ) Tokens
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16
17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32
NarrativeQA 93 75 111 51 71 75 91 108 97 154 138 143 196 126 156 201
173 133 202 200 158 128 115 86 123 109 117 76 122 120 142 206
Qasper 157 66 92 43 59 70 77 111 93 137 132 160 187 122 147 186
184 132 183 181 160 135 122 77 119 117 128 69 127 113 179 231
MultifieldQA-en 245 79 109 50 76 75 92 106 90 169 138 164 205 115 155 196
174 116 167 169 145 116 103 74 97 90 108 67 116 116 154 220
HotpotQA 145 75 115 55 73 73 92 102 88 146 128 140 190 106 140 208
172 129 194 185 146 123 109 80 105 99 125 75 127 122 173 256
2WikiMQA 198 74 110 53 69 74 91 105 89 159 133 153 196 109 143 215
178 132 181 179 139 112 96 76 102 94 120 70 118 117 161 250
MuSiQue 108 71 108 54 72 71 89 98 85 143 124 135 189 104 138 216
178 128 206 198 148 130 113 84 112 100 126 76 134 128 173 257
GovReport 332 98 113 48 73 78 88 136 125 150 176 145 142 113 136 164
212 129 148 161 145 115 79 62 103 78 94 57 99 121 138 238
QMSum 156 85 119 51 65 63 78 116 81 146 126 136 185 105 142 218
187 121 204 201 165 134 121 85 122 105 122 70 115 121 144 207
MultiNews 459 94 115 50 67 71 88 101 99 120 126 147 168 130 156 190
202 125 147 160 134 107 78 60 101 84 95 55 102 117 146 202
TREC 214 70 112 47 69 97 93 129 98 156 142 215 220 141 213 115
143 110 166 103 105 91 90 67 84 89 139 78 129 96 192 283
TriviaQA 105 62 97 52 73 89 86 106 102 178 139 178 198 140 189 199
168 118 159 139 144 103 90 66 109 102 131 71 127 124 196 256
SAMSum 126 64 103 48 71 96 93 152 118 157 130 174 173 150 189 179
168 132 177 135 135 108 98 84 101 104 120 87 114 114 168 228
PassageCount 91 72 106 48 54 58 80 111 96 119 153 143 166 147 157 225
214 147 191 203 178 106 96 72 116 116 118 70 122 130 148 243
PassageRetrieval-en 152 92 129 70 72 75 92 121 108 128 162 149 142 139 145 175
179 145 180 172 134 113 95 87 112 109 124 91 108 118 140 238
LCC 186 59 104 44 64 80 98 128 112 159 169 161 177 173 162 163
175 133 170 147 113 102 83 69 94 80 114 84 138 130 178 247
RepoBench-P 106 63 106 46 66 89 96 132 119 168 177 169 179 183 176 189
197 138 177 148 113 102 80 71 94 80 106 83 132 133 157 221
Table 7: Per-layer token processing statistics across 32 transformer layers by DynamicKV in Mistral-7B-Instruct-v0.2 model on LongBench datasets.
A.5 Task-Aware Sensitivity Study
The presented Table 7 offers a comprehensive insight into the token processing statistics per layer for 32 transformer layers using the DynamicKV method in Mistral-7B-Instruct, as applied to various datasets from LongBench. This detailed analysis reveals significant variability in how tokens are processed at different layers depending on the specific task.
In particular, our approach demonstrates that certain tasks, such as those represented by the MultiNews and GovReport datasets, exhibit exceptionally high token processing counts even in early layers, indicating a potentially greater complexity or information density in these tasks. Conversely, datasets such as TriviaQA and SAMSum show relatively lower initial processing demands, but with significant spikes in later layers, suggesting evolving complexity throughout the document or dialogue.
Moreover, the data highlights that the DynamicKV mechanism adeptly adjusts to the unique demands of each dataset, optimizing performance and resource allocation dynamically across all layers. This adaptability is crucial for handling the diverse range of tasks and input lengths encountered in real-world applications, ensuring efficient processing without compromising on understanding or accuracy.
Algorithm 1 DynamicKV in Prefill Phase
1:  Input: initial budget K/V cache list ${K}^{b}$, ${V}^{b}$, radio max $r_{max}$, update interval $m$, mean token length $wt$, window size $ws$, sequence length $S$, head dimention $hd$, input embedding of window size ${X}^{ws}\in\mathbb{R}^{ws*d}$, initial budget Attention list computed by window token and others ${A}^{b}$,
2:  Output: Compressed K/V cache ${K}^{c}$, ${V}^{c}$
3:  $bs=(wt-ws)\times r_{max}$
4:  def Update_Buffer_Length($A$, $l$):
5:       ${A}^{gather}$leftarrow cat(([${A}$ for $l$ in (1, $l$)]), 0).view(-1)
6:       $cnts$leftarrow Count_Elemnets(topk(${A}^{gather}$, k=$(wt-ws)*H*l$).indices / ($L*S$)) / $l$
7:       Compute the $norm$ of $cnts$, range in (0, 1)
8:       $BL$leftarrow [int(($bs*t$ / max($norm$))) for $t$ in $norm$]
9:       $r$leftarrow sum($BL$) / (($wt-ws$)$*L$)
10:       $BL$leftarrow [int($k/r$) for $k$ in $BL$]
11:       Return $BL$
12:  for $l\leftarrow 1\textbf{ to }L$ do
13:     Compute full KV states ${K}^{s}$, ${V}^{s}$
14:     for $h\leftarrow 1\textbf{ to }H$ do
15:        /* compute the Attention between window size token and other all token */
16:        ${A}_{l,h}$leftarrow softmax((${X}^{ws}{W}_{h}^{Q}$) $\cdot$ ${K}_{h}^{T}$).mean(dim=-2).pooling(dim=-1)
17:     end for
18:     Append ${A}_{l}$ to ${A}^{b}$ /* current ${A}_{l}$ shape is [$H$, $S$] */
19:     /* calculate current layer buffer KV cache */
20:     indices $\leftarrow$ ${A}_{l}$.topk($bs$, dim=-1).indices.unsqueeze(-1).expand(-1, -1, $hd$)
21:     ${K}_{l}^{b}$leftarrow cat((${K}^{s}$[:,:$-ws$,:].gather(dim=-2, indices),${K}^{s}$[:,$-ws$:,:]), dim=-2)
22:     ${V}_{l}^{b}$leftarrow cat((${V}^{s}$[:,:$-ws$,:].gather(dim=-2, indices),${V}^{s}$[:,$-ws$:,:]), dim=-2)
23:     /* gradually compress*/
[…]
25:        $Bl$leftarrow Update_Buffer_Length(${A}_{l}$, $l$)
26:        /* update the buffer K/V Cache*/
27:        for $i$leftarrow 1 to $l$ do
28:           ${K}_{i}^{b}$leftarrow cat((${K}_{l}^{b}$[:,:${Bl}_{i}$,:], ${K}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
29:           ${V}_{i}^{b}$leftarrow cat((${V}_{l}^{b}$[:,:${Bl}_{i}$,:], ${V}_{l}^{b}$[:,$-ws$:,:]), dim=-2)
30:        end for
31:     end if
32:  end for
33:  Update the K/V Cache ${K}^{c},{V}^{c}$ from ${K}^{b},{V}^{b}$